OpenStudy (anonymous):

5x^2m+ 40x + 75 Factor It. please Help! =)

OpenStudy (anonymous):

\[5x^2+40x+75=5(x^2+8x+15)=5(x+3)(x+5)\] i am assuming the 'm' is a typo.

OpenStudy (anonymous):

Yes! Thank You, I Didn't Catch That1 Thanks Sooo Mcuh!

OpenStudy (anonymous):

welcome;)